If p and r are prime numbers, which of the following must also be prime?
A p+r b pr+1 c p^2 + r^2 d none of above

Answer by checkley77(12844)

Answer is D) none of the above:
Proof:
3+5=8 not prime.
3*5+1=15+1=16 ditto.
3^2+5^2=9+25=34 ditto.